# Aligning Model Properties via Conformal Risk Control

### v1

@@ Appendix A Proof of multi-lambda conformal risk control
Appendix B Additional Examples
B.1 Additional monotonicity examples
@@
Table 2: Abalone, $n=4177$. Monotonically increasing on Shell_weight. $\boldsymbol{\lambda}^{\max}=(5,5)$. $g(\boldsymbol{\lambda})=\lambda^{+}+\lambda^{-}$.
$\alpha$ $\boldsymbol{\lambda}$ Metric Unconstrained Conformal Constrained
0.1 $\lambda^{+}=0.07$ MSE 5.49 5.49 5.38
$\lambda^{-}=0.00$ Risk 0.096 0.093 0
0.05 $\lambda^{+}=0.29$ MSE 5.49 5.60 5.38
$\lambda^{-}=0.47$ Risk 0.096 0.062 0
0.01 $\lambda^{+}=0.63$ MSE 5.49 6.83 5.38
$\lambda^{-}=1.95$ Risk 0.096 0.018 0
Table 3: Concrete, $n=1030$. Monotonically increasing on Cement. $\boldsymbol{\lambda}^{\max}=(2,2)$. $g(\boldsymbol{\lambda})=\lambda^{+}+\lambda^{-}$.
$\alpha$ $\boldsymbol{\lambda}$ Metric Unconstrained Conformal Constrained
0.1 $\lambda^{+}=0.004$ MSE 24.4 24.4 17.55
$\lambda^{-}=0.008$ Risk 0.081 0.081 0
0.05 $\lambda^{+}=0.092$ MSE 24.4 24.4 17.55
$\lambda^{-}=0.052$ Risk 0.081 0.075 0
0.01 $\lambda^{+}=0.710$ MSE 24.4 24.6 17.55
$\lambda^{-}=0.739$ Risk 0.081 0.043 0
B.2 Concavity results
@@
Table 4: Seoul Bike Sharing, $n=8760$. Concave on temperature. $\lambda^{\max}=100$.
@@
0.1 $\lambda=28.8$ MAE 115 117
Risk 0.31 0.090
0.05 $\lambda=45.0$ MAE 114 119
Risk 0.31 0.049
0.01 $\lambda=84.2$ MAE 115 129
Risk 0.31 0.011
Appendix C Details for a stylized examination of alignment persistence in AI models
@@
C.1 Noisy data setting
@@ C.2 RF theory
C.3 Synthetic validation under a relaxed assumption on the noise bias, and conformal correction

### v2

@@ Appendix C Proof of multi-lambda conformal risk control
Appendix D Additional Examples
D.1 Additional monotonicity examples
@@
Table 2: Abalone, $n=4177$. Monotonically increasing on Shell_weight. $\boldsymbol{\lambda}^{\max}=(5,5)$.
$\alpha$ $\boldsymbol{\lambda}$ Metric Unconstrained Adjusted Constrained
$\lambda^{+}=0.17_{\scriptsize{(\pm 0.10)}}$ MSE $5.41_{\scriptsize{(\pm 0.18)}}$ $5.46_{\scriptsize{(\pm 0.19)}}$ $5.45_{\scriptsize{(\pm 0.17)}}$
0.1 $\lambda^{-}=0.20_{\scriptsize{(\pm 0.10)}}$ Risk $0.58_{\scriptsize{(\pm 0.08)}}$ $0.09_{\scriptsize{(\pm 0.002)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
$\lambda^{+}=0.43_{\scriptsize{(\pm 0.24)}}$ MSE $5.41_{\scriptsize{(\pm 0.18)}}$ $5.63_{\scriptsize{(\pm 0.22)}}$ $5.45_{\scriptsize{(\pm 0.17)}}$
0.05 $\lambda^{-}=0.54_{\scriptsize{(\pm 0.21)}}$ Risk $0.58_{\scriptsize{(\pm 0.08)}}$ $0.04_{\scriptsize{(\pm 0.001)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
$\lambda^{+}=1.03_{\scriptsize{(\pm 0.52)}}$ MSE $5.41_{\scriptsize{(\pm 0.18)}}$ $6.67_{\scriptsize{(\pm 0.37)}}$ $5.45_{\scriptsize{(\pm 0.17)}}$
0.01 $\lambda^{-}=1.51_{\scriptsize{(\pm 0.53)}}$ Risk $0.58_{\scriptsize{(\pm 0.08)}}$ $0.01_{\scriptsize{(\pm 0.001)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
Table 3: Concrete, $n=1030$. Monotonically increasing on Cement. $\boldsymbol{\lambda}^{\max}=(2,2)$.
$\alpha$ $\boldsymbol{\lambda}$ Metric Unconstrained Adjusted Constrained
$\lambda^{+}=0.02_{\scriptsize{(\pm 0.02)}}$ MSE $24.15_{\scriptsize{(\pm 2.11)}}$ $24.18_{\scriptsize{(\pm 2.11)}}$ $22.44_{\scriptsize{(\pm 2.13)}}$
0.1 $\lambda^{-}=0.05_{\scriptsize{(\pm 0.04)}}$ Risk $0.47_{\scriptsize{(\pm 0.07)}}$ $0.07_{\scriptsize{(\pm 0.01)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
$\lambda^{+}=0.16_{\scriptsize{(\pm 0.11)}}$ MSE $24.15_{\scriptsize{(\pm 2.11)}}$ $24.37_{\scriptsize{(\pm 2.13)}}$ $22.44_{\scriptsize{(\pm 2.13)}}$
0.05 $\lambda^{-}=0.40_{\scriptsize{(\pm 0.17)}}$ Risk $0.47_{\scriptsize{(\pm 0.07)}}$ $0.04_{\scriptsize{(\pm 0.001)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
$\lambda^{+}=1.48_{\scriptsize{(\pm 0.28)}}$ MSE $24.15_{\scriptsize{(\pm 2.11)}}$ $25.45_{\scriptsize{(\pm 2.11)}}$ $22.44_{\scriptsize{(\pm 2.13)}}$
0.01 $\lambda^{-}=1.60_{\scriptsize{(\pm 0.25)}}$ Risk $0.47_{\scriptsize{(\pm 0.07)}}$ $0.004_{\scriptsize{(\pm 0.002)}}$ $0.00_{\scriptsize{(\pm 0.00)}}$
D.2 Concavity results
@@
Table 4: Seoul Bike Sharing, $n=8760$. Concave on temperature. $\boldsymbol{\lambda}^{\max}=(100,100)$
@@
$\lambda^{+}=15.93_{\scriptsize{(\pm 19.63)}}$ MAE $94.09_{\scriptsize{(\pm 2.85)}}$ $99.58_{\scriptsize{(\pm 4.03)}}$
0.1 $\lambda^{-}=24.67_{\scriptsize{(\pm 20.78)}}$ Risk $0.54_{\scriptsize{(\pm 0.02)}}$ $0.11_{\scriptsize{(\pm 0.03)}}$
$\lambda^{+}=25.40_{\scriptsize{(\pm 31.38)}}$ MAE $94.09_{\scriptsize{(\pm 2.85)}}$ $106.44_{\scriptsize{(\pm 7.67)}}$
0.05 $\lambda^{-}=38.67_{\scriptsize{(\pm 32.64)}}$ Risk $0.54_{\scriptsize{(\pm 0.02)}}$ $0.06_{\scriptsize{(\pm 0.02)}}$
$\lambda^{+}=56.93_{\scriptsize{(\pm 38.30)}}$ MAE $94.09_{\scriptsize{(\pm 2.85)}}$ $122.41_{\scriptsize{(\pm 11.41)}}$
0.01 $\lambda^{-}=71.00_{\scriptsize{(\pm 36.94)}}$ Risk $0.54_{\scriptsize{(\pm 0.02)}}$ $0.02_{\scriptsize{(\pm 0.02)}}$
Appendix E Details for a stylized examination of alignment persistence in AI models
@@
E.1 Noisy data setting
@@ E.2 RF theory
E.3 Synthetic validation under a relaxed assumption on the noise bias, and conformal correction
@@
Appendix F Connection to Mitigating LLM Hallucinations
Although it is not the main focus of this paper, we show in this section how our general methodology connects to the work of Yadkori et al. [2024]. They examine a potentially random LLM $f:{\mathcal{X}}\to\mathcal{Y}$ and assume access to a confidence score function $g:{\mathcal{X}}\to\mathbb{R}$. They use an abstention function $a:\Lambda\times{\mathcal{X}}\to\{0,1\}$ that decides whether the model should abstain from answering where $a_{\lambda}(X)=1$ if $g(X)<\lambda$ and $a_{\lambda}(X)=0$ if $g(X)\geq\lambda$ The match function $m:{\mathcal{X}}\times\mathcal{Y}\times\mathcal{Y}\to\{0,1\}$ is used to determine if a response $Y^{\prime}$ is semantically equivalent to $Y$, indicating hallucination if not.
Yadkori et al. [2024] use conformal risk control to find an optimal $\hat{\lambda}$ such that the pair $(a_{\hat{\lambda}},f)$ hallucinates at a rate less than $\alpha$. The loss function they use is $\ell:{\mathcal{X}}\times\mathcal{Y}\times\Lambda\to\mathbb{R}$ such that $\ell$ punishes failure to abstain when the response does not match the label, that is
$\ell(X,Y;\lambda)=(1-a_{\lambda}(X))(1-m(X;f(X),Y)).$ (6)
Yadkori et al. [2024] use a separate conformal procedures for setting $\lambda$ and a parameter $\beta$ with which they define $m(X;Y^{\prime},Y)=\mathbb{I}(s(X;Y^{\prime},Y)\geq\beta)$, where the similarity function $s:{\mathcal{X}}\times{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is assumed to be given. We can use our multi-lambda conformal risk control with $\boldsymbol{\lambda}=(\lambda_{a},\lambda_{m})$ to set these parameters simultaneously using a single calibration set $\{(X_{i},Y_{i})\}_{i=1}^{n}$ of ground truth query-response pairs sampled from the true data distribution ${\mathcal{D}}$. Let